APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday April 14, 2008 (MJD 54571) ---=== OBSERVING TEAM ===--- Elena Malanushenko (observer) Dan Oravetz (observer) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- The 15-degree stop became stuck in a state of "neither" according to the Interlocks during the slew from 30 degrees to 6 degrees altitude when we were opening. A light jiggle fixed the problem. See log. Elena lost her guider PGPLOT window that documents fwhm, scale, phi, etc. for no apparent reason when we were observing cartridge 3. Integrated dust count ~15,000 dust-hours. The cloud camera worked consistently for ~11 straight hours. endNight did not work, so we ran it manually and we still encountered problems with 2 of the commands. See log for details. Finished: cartridge 6, 2446 cartridge 4, 2528 Obtained signal: cartridge 3, 2876 cartridge 9, 2906 cartridge 2, 2523 ---=== OBSERVING LOG ===--- Afternoon: --------- All systems are checked and work properly. We rebooted the servers as they did not respond and rebooted photo-1 crate as id2 was unresponsive. We cleared the im pool as disc space was insufficient during two bias attempts. Spectro focus set for 5C, we did not adjust it. 00:50Z - Began ventilation. Doors and louvers opened, fans on. Night: ----- 01:38Z - Sunset. 01:40Z - When we tried slewing the telescope from stow down to 6 degrees for opening, the telescope stopped at 15.23 degrees. The Interlocks showed the 15-degree stop was in a state of "neither." In this state it had not fully retracted and upon inspection it looked like it was fully retracted, however a light jiggle of the stop resulted in its full retraction. The Interlocks confirmed that the stop had retracted. It worked fine thereafter. Perhaps this was an artifact of all of the dust we've been having lately? 01:42Z - Enclosure off. Cartridge 3, 2876, seguefaint170 ******************************** Previous signal: 54562: 54563: ----- ----- b1 r1 b2 r2 b1 r1 b2 r2 --- --- --- --- --- --- --- --- 2.6 1.5 3.2 2.0 10.1 4.3 9.9 5.8 6.2 3.2 6.9 3.8 11.1 4.8 13.1 6.5 6.6 3.7 7.0 4.5 12.1 5.1 13.3 7.4 02:10Z - gotoField. Used fk5 star and a good offset guess to acquire the field. 02:37Z - Pre-Calibs. We had yellow xsigma and wsigma values, so we adjusted sp1 and sp2 by +2000 steps each. 02:51Z - Pre-Calibs. These cals look better. Only one yellow wsigma value. 02:58Z - Science. 1500s. (S/N)^2 = 1.7 2.7 1.7 2.9 03:00Z - Elena lost her guider PGPLOT window that documents fwhm, among other things, although there didn't seem to be any obvious reason as to why it closed. She restarted SOP and the SOPGUI and it came back. 03:30Z - Science. 1500s. 1.7 2.7 2.0 3.1 03:58Z - Post-Calibs. 04:05Z - Science. 1500s. 1.9 2.3 2.0 3.2 Cartridge 6, 2446, segue139 *************************** Previous signal: 54570: ----- b1 r1 b2 r2 --- --- --- --- 8.4 5.9 2.6 3.9 4.2 3.4 1.2 3.1 4.2 2.4 1.3 3.1 04:32Z - gotoField. Used fk5 and offset guesses to acquire the field. 04:49Z - Pre-Calibs. 04:56Z - Science. 1200s. (S/N)^2 = 11.8 3.8 4.8 3.0 05:19Z - Science. 0900s. 8.8 3.1 4.5 2.5 05:37Z - Post-Calibs. This plate is DONE. Cartridge 9, 2906, segue179 *************************** 05:47Z - gotoField. Used fk5 to acquire the field. 05:53Z - Pre-Calibs. 06:00Z - Science. 0900s. (S/N)^2 = 2.3 1.5 0.8 3.2 06:18Z - Science. 1200s. 2.7 2.1 1.1 3.4 06:42Z - Science. 1500s. 3.2 2.4 1.3 3.4 07:12Z - Post-Calibs. 07:20Z - Science. 1500s. 3.5 2.6 1.4 3.6 07:49Z - Science. 1200s. 3.0 2.0 1.2 3.5 08:12Z - Post-Calibs. Cartridge 4, 2528, chunk145 *************************** 08:21Z - gotoField. Guide stars found on gotoField. 08:27Z - Pre-Calibs. 08:33Z - Science. 0900s. (S/N)^2 = 2.3 3.1 2.7 3.5 08:52Z - Science. 1500s. 5.2 5.5 6.3 6.6 09:20Z - Science. 1800s. 7.6 7.2 9.6 7.7 This plate is FINISHED. 09:45Z - Scattered clouds began moving in. Cartridge 2, 2523, chunk145 *************************** 10:03Z - gotoField. We used fk5 star and spiralPattern to acquire the field. The fk5 star was not offset significantly from the fiber (+5 offset in D resulted in a glow just off the edge of the fiber), but the presence of clouds gave made finding the star more difficult. 10:21Z - Pre-Calibs. We had to run sos_redo since the b1 channel of the flat did not reduce in SoS. We then ran it for the flat and for the first two exposures. We ran it twice on the second exposure since after the first try, the r1 channel of exposure 52014 did not reduce. 10:27Z - Science. 0900s. (S/N)^2 = 4.1 3.3 4.4 3.3 10:45Z - Science. 0900s. 5.2 3.9 5.2 3.8 11:00Z - Clouds moved out. 11:03Z - Science. 1000s. 5.8 4.7 6.4 4.2 11:38Z - Enclosure on, doors and louvers closed, fans off. 11:45Z - endNight hung on the first couple attetmps. We issued `endNight -help' and ran it manually according to the sequence of commands listed. When we got to the `verifyImagerFilesOnDisk $mjd -verbose' endNight hung and then when `createIdApoDiskLog $mjd -verbose' was issued, it hung again. We believe that these problems are related to the full pool that we discovered during afternoon checkout. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7441 00:27Z 01:30Z 100 O -80.63 -64.75 119 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 51973, plate -9999 ------- 00:02 51973 0.0 bias 00:10 51974 0.0 bias ----- sequence 51975, plate 2876 ------- 00:16 51975 10.0 flat 00:19 51976 2.0 arc 00:25 51977 3.0 arc Hartmann 00:27 51978 3.0 arc Hartmann 02:40 51979 10.0 flat 02:42 51980 2.0 arc 02:53 51981 10.0 flat 02:56 51982 2.0 arc 03:25 51983 1500.0 target ----- sequence 51984, plate 2876 ------- 03:58 51984 1501.0 target 04:01 51985 10.0 flat 04:03 51986 2.0 arc 04:32 51987 1500.0 target 04:37 51988 0.0 bias ----- sequence 51989, plate 2446 ------- 04:51 51989 10.0 flat 04:54 51990 2.0 arc 05:18 51991 1200.0 target 05:37 51992 902.0 target 05:40 51993 10.0 flat 05:42 51994 2.0 arc ----- sequence 51995, plate 2906 ------- 05:56 51995 10.0 flat 05:59 51996 2.0 arc 06:17 51997 901.0 target 06:40 51998 1201.0 target 07:09 51999 1501.0 target 07:14 52000 10.0 flat 07:16 52001 2.0 arc 07:47 52002 1500.0 target 08:11 52003 1200.0 target 08:14 52004 10.0 flat 08:17 52005 2.0 arc ----- sequence 52006, plate 2528 ------- 08:29 52006 10.0 flat 08:32 52007 2.0 arc 08:51 52008 900.0 target 09:19 52009 1502.0 target 09:52 52010 1802.0 target ----- sequence 52011, plate 2523 ------- 10:24 52011 10.0 flat 10:26 52012 2.0 arc 10:45 52013 900.0 target 11:02 52014 901.0 target 11:22 52015 1000.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:18Z 3 2876 34.86 0.0019 43.27 0.0038 209.00 0.0397 0.999980 04:10Z 3 2876 17.76 0.0016 48.13 0.0033 194.99 0.0354 0.999910 05:05Z 6 2446 160.90 0.0053 60.39 0.0053 -33.08 -0.0062 1.000080 06:05Z 9 2906 107.59 0.0033 70.39 0.0078 -85.95 0.0192 1.000200 08:36Z 4 2528 39.57 0.0033 63.74 0.0078 213.06 0.0192 0.999800 10:30Z 2 2523 -36.73 -0.0046 63.62 0.0029 149.16 0.0107 0.999950 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:18Z 3 2876 0.99998 198 1392 25 35 43.3 9.3 12 182 ---- 1.7" 04:10Z 3 2876 0.99991 889 1923 0 18 48.1 9.4 14 212 ---- 1.5" 05:05Z 6 2446 1.00008 -790 448 -50 161 60.4 9.6 14 193 ---- 1.4" 06:05Z 9 2906 1.00020 -1975 -579 -75 108 70.4 9.8 15 196 ---- 1.2" 08:36Z 4 2528 0.99990 988 1886 -75 40 63.7 9.4 11 212 ---- 1.2" 10:30Z 2 2523 0.99995 494 1552 0 -37 63.6 8.7 16 183 ---- 1.3" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:03Z 53 11 16 267 (W) 1239 - clear 00:33Z 53 12 6 258 (WSW) 1262 - clear 01:04Z 50 7 12 202 (SSW) 953 - clear 01:35Z 48 10 7 218 (SW) 1093 - clear 02:06Z 48 12 9 184 (S) 1240 - clear 02:36Z 48 11 11 195 (SSW) 1204 - clear 03:07Z 48 10 11 185 (S) 1317 - clear 03:40Z 48 10 10 189 (S) 1849 - clear 04:11Z 48 10 12 194 (SSW) 1931 - clear 04:42Z 49 11 10 197 (SSW) 2232 - clear 05:12Z 49 11 14 210 (SSW) 2115 - clear 05:43Z 49 11 13 181 (S) 2291 - clear 06:13Z 49 9 15 191 (S) 1999 - clear 06:44Z 49 10 13 185 (S) 2006 - clear 07:14Z 49 9 9 185 (S) 2014 - clear 07:44Z 48 11 12 192 (SSW) 2144 - clear 08:14Z 49 12 12 194 (SSW) 2306 - clear 08:44Z 48 10 11 201 (SSW) 2121 - clear 09:15Z 48 14 16 206 (SSW) 2230 - clear 09:45Z 47 16 12 194 (SSW) 2292 - mostly clear 10:16Z 47 18 18 204 (SSW) 2643 - scattered clouds 10:46Z 47 19 15 185 (S) 2841 - scattered clouds 11:18Z 47 21 17 214 (SW) 3308 - clear ---=== TELESCOPE STATUS ===--- 00:50Z - Began ventilation. Doors and louvers opened, fans on. 01:42Z - Enclosure off. 11:38Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 11:41Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 2 Counterweights at: 258 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 108.90 lbs, 19.41 psi Imager 119.50 lbs, 20.05 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_39 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0160 0.0200 Axial B -0.0100 0.0000 Axial C -0.0270 -0.0230 Trans D 0.0090 0.0120 Lateral E 10.0838 10.0584 Lateral F 11.6459 11.6205 GALILS Commanded: 4600. -2000. 2650. -12450. -6500. -6550. Actual: 4600. -2000. 2650. -12439. -6477. -6546. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 100.00 -200.00 PriOrient: 0.00 0.00 0.00 99.59 -200.69 SECONDARY: ---------- Focus: 0.00 Air Temp.: 10.4 Alt.: 30.074784 MIGS TONIGHT NOMINAL Axial A -0.0360 -0.0440 Axial B -0.0500 -0.0570 Axial C -0.0600 -0.0670 Trans D -0.0150 -0.0110 GALILS Commanded: 1627504. 1652946. 1470679. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.89 60.02 -4.08 -19.96 119.27 ---=== PROBLEMS IN DETAIL ===--- Collimation is set for 5C and we did not change it. sp1> Coll_motor_A -5917 01P-0000005917 Coll_motor_B -5749 02P-0000005749 Coll_motor_C -6036 01P-0000006036 Temp_Median 10.5 sp2> Coll_motor_A -301 01P-0000000301 Coll_motor_B -786 02P-0000000786 Coll_motor_C -867 01P-0000000867 Temp_Median 8.6 collimate, 51977 Reading /data/spectro/54571/sdR-r1-00051977.fit Reading /data/spectro/54571/sdR-r1-00051978.fit COLLIMATE: Min offset = -1.04 pix COLLIMATE: Max offset = -0.10 pix COLLIMATE: RMS across CCD = 0.28 pix COLLIMATE: Mean offset = -0.63 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 5782 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 5782" Log file = Collimate-54571-r1-00051977.log PostScript file = Collimate-54571-r1-00051977.ps Reading /data/spectro/54571/sdR-r2-00051977.fit Reading /data/spectro/54571/sdR-r2-00051978.fit COLLIMATE: Min offset = -1.08 pix COLLIMATE: Max offset = -0.06 pix COLLIMATE: RMS across CCD = 0.29 pix COLLIMATE: Mean offset = -0.40 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 3691 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 3691" Log file = Collimate-54571-r2-00051977.log PostScript file = Collimate-54571-r2-00051977.ps % Program caused arithmetic error: Floating illegal operand